40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 326     Entry time: Wed Feb 20 16:24:37 2008
Author: steve 
Type: Update 
Category: PSL 
Subject: the laser is recovering slowly 
Head temp is still 20.5C and decreasing slowly.
Power output 2.9W
NPRO power 22 mW is increasing as head is cooling down
Attachment 1: laserecoverring.jpg  89 kB  | Hide | Hide all
laserecoverring.jpg
ELOG V3.1.3-